# SPECIAL RELATIVITY

### v11

@@ 12.1 Pseudo-Gravitation Fields and the Langevin Metric
Suppose that the reference frame $S$ is at rest with respect to the fixed stars. Let $(ct,x,y,z)$ represent the Lorentz coordinates in the inertial frame $S$, while $S_{n}$ is a system initially at rest relative to $S$ that undergoes uniform acceleration along the $x$-axis, reaching a velocity $v$. The inverse Galilean transformation $x\to x_{n}+vt$, $t\to t_{n}$ defines a new reference frame, $S_{n}$, which is uniformly accelerated in the $x$-direction, attaining a velocity $v$ relative to $S$. After the acceleration, the frame $S$ is, reciprocally, moving along the $x_{n}$-axis with velocity $-v$ relative to $S_{n}$.
When the system $S_{n}$ begins moving with constant velocity, it becomes an inertial frame of reference. At first glance, there appears to be symmetry (or reciprocity) between the inertial frame $S$ and the inertial frame $S_{n}$. According to the principle of relativity (also known as the principle of reciprocity), an observer accelerated relative to the fixed stars should measure the Minkowski metric as $ds^{2}=c^{2}dt_{n}^{2}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$. This metric is reciprocal to the one measured by the inertial observer: $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$. So, where does the asymmetry arise? The analysis of the equivalence principle reveals that no additional complexities are required to explain the asymmetry between an inertial frame and an accelerated frame (relative to the fixed stars). The resolution of this paradox lies in identifying a pseudo-gravitational potential within the system $S_{n}$, which acts as the source of the asymmetry.
@@
To minimize the mathematical complexity of the discussion, we will restrict our analysis to terms up to second order in $v/c$. Consider a reference frame $S$, in which each point undergoes constant acceleration $g=-|g|$ in the negative $x_{n}$-direction for a duration $T=v/|g|$. In this setup, the system $S$ is effectively in free fall within a pseudo-gravitational field. It becomes evident that the pseudo-gravitational acceleration is given by the gradient of a scalar potential, i.e., $g=-\partial\phi/\partial x_{n}$. A straightforward calculation shows that at time $t=T$, the coordinate transforms as $x_{n}=gT^{2}/2+x$ at $t=T$. Once the system $S$ transitions to uniform motion (constant velocity), the pseudo-gravitational field vanishes. To determine the pseudo-gravitational potential difference between the two frames, we compute $\phi=(\partial\phi/\partial x_{n})gT^{2}/2=-v^{2}/2$. Here, we have assumed that $\phi=0$ at $x_{n}=x$.
@@
Let us first examine the behavior of a clock in free fall. From the perspective of an observer located at $\phi_{2}$, the clock in the freely falling frame appears to tick faster, since it remains associated with the higher potential $\phi_{1}$, regardless of its instantaneous position. Radiation from a standard source within the free-falling frame will therefore appear blue-shifted relative to radiation from a standard source at rest at $\phi_{2}$ 434343 To understand the gravitational blue-shift of a light source in free fall, one must analyze the source’s emission mechanism and its behavior while in motion. The frequency of dipole radiation is proportional to the electron mass. In a gravitational field, mass varies with gravitational potential, according to the relation $m=m_{0}/\sqrt{1+2\phi/c^{2}}$. As a result, the electron mass is greater at points with lower gravitational potential..
@@
We now have all the quantities we set out to determine. Let us bring them together into the expression for the new space-time variables. The new independent variables $(x_{n}^{\mathrm{(p)}},t_{n}^{\mathrm{(p)}})$ can be expressed in terms of the original independent variables $(x_{n},t_{n})$ as follows: $ct_{n}^{\mathrm{(p)}}=t_{n}[1-v^{2}/(2c^{2})]-(vx_{n}/c^{2})[1+v^{2}/(2c^{2})]$, $x_{n}^{\mathrm{(p)}}=x_{n}[1+v^{2}/(2c^{2})]$. Above we made a simplification in calculating time and space transformations by considering only low velocities. Despite this approximation, the physical principles underlying the space-time transformation were clearly illustrated. The next natural question is: What is the form of the transformation for arbitrary velocity? The answer is straightforward: $t_{n}^{\mathrm{(p)}}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x_{n}^{\mathrm{(p)}}=x_{n}/\sqrt{1-v^{2}/c^{2}}$. The quantity $dt_{n}^{\mathrm{(p)}}$ represents physical time, which is independent on the choice of coordinate time. It is important to note that $t_{n}$ in the expression for $t_{n}^{\mathrm{(p)}}$ corresponds to the coordinate time in the accelerated frame. Physical time governs the actual flow of events in physical processes. Likewise, the quantity $|dx_{n}^{\mathrm{(p)}}|$ represents the physical distance between two points in three-dimensional space. The space-time interval, when expressed in terms of physically measurable differentials, becomes a diagonal quantity. These measurable differentials, $dt_{n}^{\mathrm{(p)}}$ and $dx_{n}^{\mathrm{(p)}}$, are linear combinations of the coordinate time differential $dt_{n}$ and the spatial differential $dx_{n}$, with coefficients determined by components of the metric tensor. In the new physical coordinates, the space-time interval takes the form: $ds^{2}=(cdt_{n}^{\mathrm{(p)}})^{2}$ $(dx_{n}^{\mathrm{(p)}})^{2}$ $=(1-v^{2}/c^{2})c^{2}dt^{2}_{n}-2vdx_{n}dt_{n}-dx_{n}^{2}$.
@@ 12.2 The Equivalence Principle and Gravitational Physics
We should make one further remark regarding the equivalence principle. [^44] For a general discussion of the equivalence principle, we recommend reading [50]. It is a common belief that the equivalence principle forms the foundation of general relativity. However, this view is not entirely accurate. The gravitational field is not merely a universal force acting identically on all bodies. According to Einstein’s insight, gravity is intimately connected to the structure of space-time itself. To grasp this idea, we must first understand the role of the Riemann curvature tensor, which, in a precise sense, is the gravitational field. Only when this tensor vanishes can we say that no gravitational field exists. This observation alone suffices to demonstrate that the equivalence principle, as traditionally formulated, does not universally apply. The equivalence principle holds for mechanical processes but fails when applied to electrodynamical phenomena. This distinction implies that it is indeed possible to determine, through internal measurements, whether a system is in an inertial frame or in free fall within a homogeneous gravitational field. For example, in such a field, a free electric charge undergoes uniform acceleration relative to an inertial frame and, as a result, emits radiation. This radiation occurs regardless of whether the field is strong or weak. Conversely, in a uniformly accelerated frame (absent gravity), the free charge remains unaccelerated relative to the inertial frame and, according to classical electrodynamics, does not radiate. Hence, the equivalence principle is explicitly violated in this context.
@@ 12.3 Time Dilation
It is commonly believed that special relativity is a reciprocal theory. However, an instructive counterexample to this reciprocity is provided by the time dilation effect. Consider the passage of time in two inertial reference frames. One frame, denoted by $S$, is considered to be at rest relative to the fixed stars. The second frame, $S_{n}$, undergoes acceleration relative to $S$ until it reaches a velocity $v$ along the $x$-axis. As previously discussed, the transformation from an inertial to an accelerated frame was apparently first introduced by Langevin. A central mathematical approach in the analysis of time dilation is the smooth matching—or tailoring—of the metric tensor. The transformation to the accelerated frame, with coordinates denoted by subscript ”n”, takes a Galilean form: $t_{n}=t$, $x_{n}=x-vt$ Following this transformation, the metric in the inertial frame transitions continuously into the Langevin metric of the accelerated inertial frame $S_{n}$ (see Chapter 5 for details).
@@
Let us now consider a different scenario. Suppose that frame $S_{n}$ accelerates relative to the inertial frame $S$ until it reaches a velocity $v$. Then, two new frames, $S_{n}^{(1)}$ and $S_{n}^{(2)}$, each accelerate relative to $S_{n}$ to the same speed $v$, but in opposite directions. From the perspective of $S_{n}$, this setup is not symmetric, and the resulting time dilation experienced in the accelerated frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ differs. In frame $S_{n}$, frame $S_{n}^{(1)}$ moves with velocity $-v$, so the proper time in this frame is simply $d\tau_{n}^{(1)}=dt=dt_{n}$. Meanwhile, frame $S_{n}^{(2)}$ moves with velocity $v$ relative to $S_{n}$. Using the Langevin metric, $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}$, we find that the proper time in frame $S_{n}^{(2)}$ is given by $d\tau_{n}^{(2)}=\sqrt{1-4v^{2}/c^{2}}dt_{n}$. This analysis has been carried out from the point of view of an observer at rest in frame $S_{n}$. It is natural to ask: how would the situation appear to an observer in the inertial frame $S$? From the perspective of frame $S$ the frame $S_{n}^{(2)}$ moves with velocity $dx/dt=2v$. Applying the Minkowski metric, we again obtain: $d\tau_{n}^{(2)}=\sqrt{1-4v^{2}/c^{2}}dt$, which is consistent with our previous result. It is worth noting that the mathematics involved in relating proper times is remarkably straightforward in this case. This simplicity arises because we are using Galilean boosts and the Galilean addition of velocities.
@@ 12.4 The Light-Clock and Observations by a Non-Inertial Observer
To understand how a clock slows down in an accelerated system, we need to examine the internal workings of the clock and observe what happens as it moves. We’ll explore this effect using a simplified example—a device known as a light clock. This clock consists of a meter stick with mirrors at both ends; a ”tick” corresponds to a pulse of light bouncing between the mirrors. Now, let us consider two configurations of the light clock: one with its length parallel to the direction of motion and the other oriented orthogonally. According to the principle of relativity, all clocks—regardless of their construction—must measure time consistently in all inertial frames. We will demonstrate that, regardless of the orientation of the light clock relative to its motion, the elapsed time measured remains the same. Two light clocks illustrating these configurations are shown in Fig. 47.
@@
The spatial geometry in an accelerated frame is described by the differential spatial line element defined within it. While spatial geometry possesses a degree of conventionality due to coordinate dependence, the proper spatial line element, denoted $dl_{n}$, holds particular significance, as it defines a coordinate-independent spatial geometry in the accelerated frame. The relation $dl_{n}^{2}=(-g_{\alpha\beta}+g_{0\alpha}g_{0\beta}/g_{00})dx_{\alpha}dx_{\beta}$ establishes the connection between the three-dimensional spatial line element and the four-dimensional spacetime metric. It can be shown that $dl_{n}$ remains invariant under coordinate transformations within the same reference frame. In contrast, the spatial line element $dl$, obtained by setting $dt_{n}=0$ in Eq.(15) with $dl^{2}=dx_{n}^{2}$, differs from $dl_{n}$.
Using Eq. (15), we obtain the relation $dl_{n}^{2}=dx_{n}^{2}/(1-v^{2}/c^{2})$, which defines the spatial geometry in the accelerated frame. This equation is asymmetric with respect to the lengths $dx_{n}=dl$ and $dl_{n}$, as it connects the physical length $dl$ of a rod measured in the inertial frame $S$ with the physical length $dl_{n}$ in the accelerated frame $S_{n}$. Our analysis, based on the use of standard measuring rods within the accelerated frame to probe its geometrical properties, shows that a coordinate distance $dx_{n}$ corresponds to a physical length of $dx_{n}/\sqrt{1-v^{2}/c^{2}}$. Thus, we can directly determine the asymmetry in the length of the meter stick from the metric. This can be interpreted as an effect of acceleration with respect to the fixed stars, suggesting that space in an accelerated frame is inherently non-Euclidean.
It is important to emphasize that the Langevin metric characterizes the measurements of an accelerated observer as interpreted by an inertial observer who employs the simplest form of clock synchronization—namely, maintaining the same set of synchronized clocks used when the light clock was at rest. This method preserves Einstein synchronization, established via light signals emitted from a stationary dipole source. Within this framework of absolute simultaneity, the inertial observer perceives the accelerated observer as measuring the speed of light along the $x$-axis as $c-v$ in the positive direction and $-c-v$ in the negative direction. Furthermore, the physical meter stick (with mirrors mounted on it), as measured by the accelerated observer, appears compressed along the direction of motion relative to the fixed stars.

### v12

@@ 12.1 Pseudo-Gravitation Fields and the Langevin Metric
Suppose that the reference frame $S$ is at rest with respect to the fixed stars. Let $(ct,x,y,z)$ represent the Lorentz coordinates in the inertial frame $S$, while $S_{n}$ is a system initially at rest relative to $S$ that undergoes uniform acceleration along the $x$-axis, reaching a velocity $v$. The inverse Galilean transformation $x\to x_{n}+vt$, $t\to t_{n}$ defines a new reference frame, $S_{n}$, which is uniformly accelerated in the $x$-direction, attaining a velocity $v$ relative to $S$. After the acceleration, the frame $S$ is, reciprocally, moving along the $x_{n}$-axis with velocity v-v relative to $S_{n}$.
When the system $S_{n}$ begins moving with constant velocity, it becomes an inertial frame of reference. At first glance, there appears to be symmetry (or reciprocity) between the inertial frame $S$ and the inertial frame $S_{n}$. According to the principle of relativity (also known as the principle of reciprocity), an observer accelerated relative to the fixed stars should measure the Minkowski metric as $ds^{2}=c^{2}dt_{n}^{2}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$. This metric is reciprocal to the one measured by the inertial observer: $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$. So, where does the asymmetry arise? The analysis of the equivalence principle reveals that no additional complexities are required to explain the asymmetry between an inertial frame and an accelerated frame. The resolution of this paradox lies in identifying a pseudo-gravitational potential within the system $S_{n}$, which acts as the source of the asymmetry.
@@
To minimize the mathematical complexity of the discussion, we will restrict our analysis to terms up to second order in $v/c$. Consider a reference frame $S$, in which each point undergoes constant acceleration g=|g|g=-|g| in the negative $x_{n}$-direction over a time interval $T=v/|g|$. In this configuration, frame $S$ is effectively in free fall within a pseudo-gravitational field experienced in the accelerating frame $S_{n}$. It becomes evident that the pseudo-gravitational acceleration is given by the gradient of a scalar potential, i.e., $g=-\partial\phi/\partial x_{n}$. A straightforward calculation shows that at time $t=T$, the coordinate transforms as $x_{n}=gT^{2}/2+x$ at $t=T$. Once the system $S$ transitions to uniform motion (i.e. constant velocity), the pseudo-gravitational field vanishes. To determine the pseudo-gravitational potential difference between the two frames, we compute $\phi=(\partial\phi/\partial x_{n})gT^{2}/2=-v^{2}/2$, where we have chosen the potential to be zero at $x_{n}=x$, i.e., $\phi=0$ at that point.
@@
Let us first examine the behavior of a clock in free fall. From the perspective of an observer located at $\phi_{2}$, the clock in the freely falling frame appears to tick faster, since it remains associated with the higher potential $\phi_{1}$, regardless of its instantaneous position. Radiation from a standard source within the free-falling frame will therefore appear blue-shifted relative to radiation from a standard source at rest at $\phi_{2}$. 444444 To understand the gravitational blue-shift of a light source in free fall, one must analyze the source’s emission mechanism and its behavior while in motion. The frequency of dipole radiation is proportional to the electron mass. In a gravitational field, mass varies with gravitational potential, according to the relation $m=m_{0}/\sqrt{1+2\phi/c^{2}}$. As a result, the electron mass is greater at points with lower gravitational potential.
@@
We now have all the quantities we set out to determine. Let us bring them together into the expression for the new space-time variables. The new independent variables $(x_{n}^{\mathrm{(p)}},t_{n}^{\mathrm{(p)}})$ can be expressed in terms of the original independent variables $(x_{n},t_{n})$ as follows: $ct_{n}^{\mathrm{(p)}}=t_{n}[1-v^{2}/(2c^{2})]-(vx_{n}/c^{2})[1+v^{2}/(2c^{2})]$, $x_{n}^{\mathrm{(p)}}=x_{n}[1+v^{2}/(2c^{2})]$. Above we made a simplification in calculating time and space transformations by considering only low velocities. Despite this approximation, the physical principles underlying the space-time transformation were clearly illustrated. The next natural question is: What is the form of the transformation for arbitrary velocity? The answer is straightforward: $t_{n}^{\mathrm{(p)}}=t_{n}\sqrt{1-v^{2}/c^{2}}-(vx_{n}/c^{2})/\sqrt{1-v^{2}/c^{2}}$, $x_{n}^{\mathrm{(p)}}=x_{n}/\sqrt{1-v^{2}/c^{2}}$. The quantity $dt_{n}^{\mathrm{(p)}}$ represents physical time, which is independent on the choice of coordinate time. It is important to note that $t_{n}$ in the expression for $t_{n}^{\mathrm{(p)}}$ corresponds to the coordinate time in the accelerated frame. Physical time governs the actual flow of events in physical processes. Likewise, the quantity $|dx_{n}^{\mathrm{(p)}}|$ represents the physical distance between two points in three-dimensional space. The space-time interval, when expressed in terms of physically measurable differentials, becomes a diagonal quantity. These measurable differentials, $dt_{n}^{\mathrm{(p)}}$ and $dx_{n}^{\mathrm{(p)}}$, are linear combinations of the coordinate time differential $dt_{n}$ and the spatial differential $dx_{n}$, with coefficients determined by components of the metric tensor. In the new physical coordinates, the space-time interval takes the form: $ds^{2}=(cdt_{n}^{\mathrm{(p)}})^{2}$ - $(dx_{n}^{\mathrm{(p)}})^{2}$ $(1-v^{2}/c^{2})c^{2}dt^{2}_{n}-2vdx_{n}dt_{n}-dx_{n}^{2}$.
@@ 12.2 The Equivalence Principle and Gravitational Physics
We should add a further remark concerning the equivalence principle. [^45] For a general discussion of the equivalence principle, we recommend [50]. It is commonly believed that the equivalence principle lies at the foundation of general relativity. However, this perspective is not entirely accurate. Gravity, as understood through Einstein’s insight, is not merely a universal force acting identically on all bodies—it is fundamentally tied to the geometry of space-time itself.
To appreciate this idea, one must consider the role of the Riemann curvature tensor, which, in a precise mathematical sense, represents the gravitational field. The vanishing of this tensor is the necessary and sufficient condition for the absence of gravity. This fact alone shows that the traditional formulation of the equivalence principle has only limited applicability.
While the equivalence principle holds for mechanical processes, it breaks down when applied to electrodynamical phenomena. This distinction implies that internal measurements can, in principle, distinguish between an inertial frame and a frame in free fall within a homogeneous gravitational field.
For instance, consider a free electric charge in such a gravitational field. Relative to an inertial frame, it experiences uniform acceleration and consequently emits radiation—regardless of whether the gravitational field is weak or strong. In contrast, in a uniformly accelerated frame in flat space-time (i.e., in the absence of gravity), a freely falling charge does not accelerate relative to the inertial frame and, according to classical electrodynamics, does not radiate. Thus, in this context, the equivalence principle is explicitly violated.
@@ 12.3 Time Dilation
It is commonly believed that special relativity is a reciprocal theory. However, an instructive counterexample to this reciprocity is provided by the time dilation effect. Consider the passage of time in two inertial reference frames. One frame, denoted by $S$, is considered to be at rest relative to the fixed stars. The second frame, $S_{n}$, undergoes acceleration relative to $S$ until it reaches a velocity $v$ along the $x$-axis. As previously discussed, the transformation from an inertial to an accelerated frame was apparently first introduced by Langevin. A central mathematical approach in the analysis of time dilation is the smooth matching—or tailoring—of the metric tensor. The transformation to the accelerated frame, with coordinates denoted by subscript ”n”, takes a Galilean form: $t_{n}=t$, $x_{n}=x-vt$ . Following this transformation, the metric in the inertial frame transitions continuously into the Langevin metric of the accelerated inertial frame $S_{n}$ (see Chapter 5 for details).
@@
Let us now consider a different scenario. Suppose that frame $S_{n}$ accelerates relative to the inertial frame $S$ until it reaches a velocity $v$. Then, two new frames, $S_{n}^{(1)}$ and $S_{n}^{(2)}$, each accelerate relative to $S_{n}$ to the same speed $v$, but in opposite directions. From the perspective of $S_{n}$, this setup is not symmetric, and the resulting time dilation experienced in the accelerated frames $S_{n}^{(1)}$ and $S_{n}^{(2)}$ differs. In frame $S_{n}$, frame $S_{n}^{(1)}$ moves with velocity v-v, so the proper time in this frame is simply $d\tau_{n}^{(1)}=dt=dt_{n}$. Meanwhile, frame $S_{n}^{(2)}$ moves with velocity $v$ relative to $S_{n}$. Using the Langevin metric, $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}$, we find that the proper time in frame $S_{n}^{(2)}$ is given by $d\tau_{n}^{(2)}=\sqrt{1-4v^{2}/c^{2}}dt_{n}$. This analysis has been carried out from the point of view of an observer at rest in frame $S_{n}$. It is natural to ask: how would the situation appear to an observer in the inertial frame $S$? From the perspective of frame $S$ the frame $S_{n}^{(2)}$ moves with velocity $dx/dt=2v$. Applying the Minkowski metric, we again obtain: $d\tau_{n}^{(2)}=\sqrt{1-4v^{2}/c^{2}}dt$, which is consistent with our previous result. It is worth noting that the mathematics involved in relating proper times is remarkably straightforward in this case. This simplicity arises because we are using Galilean boosts and the Galilean addition of velocities.
@@ 12.4 The Light-Clock and Observations by a Non-Inertial Observer
To understand how a clock slows down in an accelerated system, we need to examine the internal workings of the clock and observe what happens as it moves. We’ll explore this effect using a simplified example—a device known as a light clock. This clock consists of a meter stick with mirrors at both ends; a ”tick” corresponds to a pulse of light bouncing between the mirrors. Now, let us consider two configurations of the light clock: one with its length parallel to the direction of motion and the other oriented orthogonally. According to the principle of relativity (i.e. pseudo-Euclidean spacetime geometry), all clocks—regardless of their construction—must measure time consistently in all inertial frames. We will demonstrate that, regardless of the orientation of the light clock relative to its motion, the elapsed time measured remains the same. Two light clocks illustrating these configurations are shown in Fig. 47.
@@
The spatial geometry in an accelerated frame is characterized by the differential spatial line element defined within that frame. Although spatial geometry is inherently coordinate-dependent and thus exhibits a degree of conventionality, the proper spatial line element, denoted by $dl_{n}$, holds particular significance. It defines a coordinate-independent notion of spatial distance within the accelerated frame. By contrast, the spatial line element $dl$, obtained by setting $dt_{n}=0$ in Eq. (15) such that $dl^{2}=dx_{n}^{2}$, differs from $dl_{n}$.
Now, how do physical notions of length and time manifest in the coordinate system $(t_{n},x_{n},y_{n},z_{n})$? In an accelerated frame, consider the spacetime interval given by $ds^{2}=c^{2}g_{00}dt^{2}+2cg_{01}dxdt+g_{11}dx^{2}$. We can decompose this expression as follows: $ds^{2}=c^{2}[\sqrt{g_{00}}dt+g_{01}dx/(c\sqrt{g_{00}})]^{2}+(g_{11}-g_{01}^{2}/g_{00})dx^{2}$. Here, the quantity $dT^{2}=[\sqrt{g_{00}}dt+g_{01}dx/(c\sqrt{g_{00}})]^{2}$ has a time-like character and represents the square of physical time in the accelerated frame, while $dX^{2}=(g_{11}-g_{01}^{2}/g_{00})dx^{2}$ is space-like and corresponds to the square of physical length. These quantities are directly tied to the spacetime interval in a manner analogous to the Minkowski metric in special relativity [16]. They are thus physically meaningful and invariant under coordinate transformations within the accelerated frame.
The relation $dl_{n}^{2}=(-g_{11}+g_{01}g_{01}/g_{00})dx^{2}$ establishes the connection between the three-dimensional spatial line element and the four-dimensional spacetime metric. Using Eq. (15), we obtain the relation $dl_{n}^{2}=dx_{n}^{2}/(1-v^{2}/c^{2})$, which defines the spatial geometry in the accelerated frame. This equation is asymmetric with respect to the lengths $dx_{n}=dl$ and $dl_{n}$, as it connects the physical length $dl$ of a rod measured in the inertial frame $S$ with the physical length $dl_{n}$ in the accelerated frame $S_{n}$. Our analysis, based on the use of standard measuring rods within the accelerated frame to probe its geometrical properties, shows that a coordinate distance $dx_{n}$ corresponds to a physical length of $dx_{n}/\sqrt{1-v^{2}/c^{2}}$. Thus, we can directly determine the asymmetry in the length of the meter stick from the metric. This can be interpreted as an effect of acceleration with respect to the fixed stars, suggesting that space in an accelerated frame is inherently non-Euclidean.
It is important to emphasize that the Langevin metric characterizes the measurements of an accelerated observer as interpreted by an inertial observer who employs the simplest form of clock synchronization—namely, maintaining the same set of synchronized clocks used when the light clock was at rest. This method preserves Einstein synchronization, established via light signals emitted from a stationary dipole source. Within this framework of absolute simultaneity, the inertial observer perceives the accelerated observer as measuring the speed of light along the $x$-axis as cvc-v in the positive direction and cv-c-v in the negative direction. Furthermore, the physical meter stick (with mirrors mounted on it), as measured by the accelerated observer, appears compressed along the direction of motion relative to the fixed stars.